Olivia can type a report in 2 hours and john takes 3 hours. If the two of them work together, how long will it take to type up the report
Answer by lwsshak3(11628)

Olivia can type a report in 2 hours and john takes 3 hours. If the two of them work together, how long will it take to type up the report
***
1/2=Olivia's work rate
1/3=john's work rate
1/x=work rate working together
..
sum of indv work rates=work rate working together
1/2+1/3=1/x
5/6=1/x
x=6/5=1.2
If the two of them work together, how long will it take to type up the report? 1.2 hrs